# Kapranov degrees

### v3

@@ 1. Introduction
For a finite set $S$ with cardinality $|S|\geq 3$, let $\overline{M}_{0,S}$ be the Deligne–Mumford–Knudsen moduli space of stable rational curves with marked points labelled by $S$. We write $\overline{M}_{0,n}$ for $\overline{M}_{0,[n]}$, where $[n]\coloneqq\{1,\dots,n\}$ for an integer $n\geq 3$. For a pair $(S,i)$ consisting of a subset $S\subseteq[n]$ with $|S|\geq 3$ and an element $i\in S$, let
$f_{S}$ $\colon\overline{M}_{0,n}\to\overline{M}_{0,S}\ \text{ be the forgetful map induced by forgetting the markings not in $S$, and}$
$|\psi_{i}|$ $\colon\overline{M}_{0,S}\to\mathbb{P}^{|S|-3}\ \text{ be the \emph{Kapranov map} induced by the complete linear series of the \emph{psi-class} $\psi_{i}$},$
@@
For a collection $(S_{1},i_{1}),\dots,(S_{n-3},i_{n-3})$ of $n-3$ such pairs, we define its Kapranov degree to be the intersection number
@@
$\overline{M}_{0,n}\hookrightarrow\prod_{\begin{subarray}{c}(S,i)\text{ such that}\\ i\in S\subseteq[n]\text{ and }|S|\geq 3\end{subarray}}\mathbb{P}^{|S|-3}$
@@
    When $|S_{j}|=4$ for all $j$, Kapranov degrees in this case were studied by Silversmith [Silversmith] under the name of cross-ratio degrees, because the divisor class $X_{S,i}$ when $|S|=4$ is the pullback of the hyperplane class under the map $\overline{M}_{0,n}\to\mathbb{P}^{1}$ given by taking the cross-ratio of the four points marked by $S$. That is, the Kapranov degree in this case can be interpreted as the number of ways to choose $n$ marked points on $\mathbb{P}^{1}$ with $n-3$ prescribed cross-ratios.
@@
    When further $n=2v$, and the size 4 subsets $S_{j}$ are obtained from a graph $G$ with vertex set $[v]$ and $32v-$ 3 edges by assigning to each edge $(a,b)$ the subset $\{a,b,a+v,b+v\}$, the cross-ratio degree counts the number of (complex) realizations of $G$ on the $2$-dimensional sphere up to isometries for any fixed generic edge lengths of $G$ [LamanSphere].
@@
    When $i_{j}=n$ for all $j$, the Kapranov degrees were computed in [CTHypertrees, Theorem 3.2]. In this case the Kapranov degree is either 00 or $1$.
@@
$\operatorname{res}_{P}(S,i)=\begin{cases}X_{S,i}&\text{if $S\subseteq P$}\\ X_{(S\setminus i)\cup\star,\star}&\text{if $S\cap P=S\setminus i$}\\ X_{(S\cap P)\cup\star,i}&\text{if $S\cap P\supseteq\{i,j\}$ for some $i\neq j\in S$}\\ 1&\text{otherwise, i.e.,\ if $S\cap P=\{i\}$ or $S\cap P\subseteq S\setminus\{i,j\}$ for some $i\neq j\in S$.}\end{cases}$
One may find an illustration of these cases in Figure 1 helpful for visualizing the above conditions. We similarly define $\operatorname{res}_{Q}(S,i)$ by replacing $P$ with $Q$ in the above. It follows from the definition that $\operatorname{res}_{P}(S,i)=1$ if and only if $\operatorname{res}_{Q}(S,i)\neq 1$ and vice versa.
@@
When $|S_{j}|=4$ for all $j\in[n-3]$, i.e., the case of cross-ratio degrees, B was proven in [LamanSphere]. It was also proven in [Goldner, Lemma 3.11] using tropical geometry (see [Silversmith]). We prove Theorem B in Section 3 by writing $X_{S,i}-X_{S\setminus a,i}$ as a sum of boundary divisors and using that the boundary divisors are products of smaller $\overline{M}_{0,n}$’s. Theorem B gives a useful algorithm for computing Kapranov degrees. [^1] An implementation in SageMath is available at https://github.com/MattLarson2399/KapranovDegrees. For the case of cross-ratio degrees, the recursion further admits a combinatorial interpretation in terms of trivalent trees (Corollary 3.3).
@@
Lastly, we give an explicit formula for the special case of Kapranov degrees where all the $S_{j}$ share a subset of size $3$, say $\{n-2,n-1,n\}$, and each $i_{j}$ lies in $\{n-2,n-1,n\}$. For a pair $(T,j)$ consisting of $T\subseteq[n-3]$ and $j\in\{n-2,n-1,n\}$, let $Y_{T,j}=X_{T\cup\{n-2,n-1,n\},j}$. For $n-3$ such pairs $(T_{1},j_{1}),\dotsc,(T_{n-3},j_{n-3})$, we define a $3$-transversal to be a map $t\colon[n-3]\to\{n-2,n-1,n\}$ such that there is a bijection $m\colon[n-3]\to[n-3]$ satisfying $m(i)\in T_{i}$ and $t(m(i))=j_{i}$ for all $i\in[n-3]$.
@@ 2.1. Reduction to the case of cross-ratio degrees
We will apply the Hall–Rado theorem to the matroid known as the third Dilworth truncation of the boolean matroid, denoted $D_{3}(U_{n,n})$. See, for example, [Whittle]. Its ground set is the collection $\binom{[n]}{4}$ of all subsets of $[n]$ of size $4$, and its set of bases consists of collections of $n-3$ subsets of size $4$ that satisfy the Cerberus condition. In [Athanasiadis, Theorem 3.2], the flats of $D_{3}(U_{n,n})$ are calculated to be as follows. They are in bijection with anti-chains $\{T_{1},\dotsc,T_{k}\}$ in the poset of subsets of $[n]$ that satisfy the conditions
@@
For the claim, we first note that the closure of $T_{S,i}$ in $D_{3}(U_{n,n})$ is the flat $F_{S}$ corresponding to $\{S\}$ because $F_{S}$ contains $T_{S,i}$ and they both have rank $|S|-3$. We now need show that the collection of flats $F_{S_{1}},\dotsc,F_{S_{n-3}}$ of $D_{3}(U_{n,n})$ satisfies (Hall–Rado). Suppose on the contrary that it fails (Hall–Rado), and take a minimal $J\subseteq[n-3]$ witnessing the failure. For each $j\in J$, because the flat $F_{S_{j}}$ is connected, it is contained in a connected component of the join $\bigvee_{j\in J}F_{S_{j}}$. Since ranks add over connected components, the minimality of $J$ implies that $\bigvee_{j\in J}F_{S_{j}}$ is a connected flat of rank less than $|J|$ in $D_{3}(U_{n,n})$. But the smallest connected flat containing $F_{S_{j}}$ for all $j\in J$ is the flat $F_{\bigcup_{j\in J}S_{j}}$, which has rank $|\bigcup_{j\in J}S_{j}|-3\geq|J|$ because of the Cerberus condition. ∎
@@ 2.2. Positivity for cross-ratio degrees
It suffices to show that the $(n-3)\times n$ matrix whose rows are given by the $d\operatorname{cr}_{S_{i}}$ has full rank; we may check this after scaling each row to clear denominators. By the computation above, if $S_{j}=\{j_{1},j_{2},j_{3},j_{4}\}$, then $i$-th row has $(-1)^{\ell+1}\prod_{a<b\in S_{i},a,b\not=j_{\ell}}(x_{a}-x_{b})$ in the $j_{\ell}$-th entry and is zero otherwise.
Let $M$ be the $(n-3)\times n$ matrix whose $i$-th column is given by $(1,x_{i},\dotsc,x_{i}^{n-4})$. We apply Theorem 2.5 with $m=n-3$ and each $T_{i}=S_{i}$. Because $S_{1},\dotsc,S_{n-3}$ satisfies the Cerberus condition, there is an invertible $(n-3)\times(n-3)$ matrix $G$ such $(GM)_{ij}=0$ unless $j\in S_{i}$. By Proposition 2.6, we may pick $G$ such that $(GM)_{ij}=\prod_{k\in[n]\setminus S_{i}}(x_{j}-x_{k})$.
@@
Because the GM-MDS theorem holds over a field of any characteristic, Proposition 2.4 implies that, in any characteristic, the rational functions corresponding to cross-ratios form a separating transcendence basis when the Cerberus condition is satisfied. This implies that the map $\overline{M}_{0,n}\to(\mathbb{P}^{1})^{n-3}$ given by $n-3$ cross-ratios is generically étale when it is dominant.
@@
Let $I$ be the group of Euclidean isometries of the plane $\mathbb{R}^{2}$. For a graph $G$ with vertex set $[v]$ and edges $E$, consider the map $\Psi_{G}\colon(\mathbb{R}^{2})^{[v]}/I\to\mathbb{R}^{E}$ given by $[p\colon[v]\to\mathbb{R}^{2}]\mapsto\big{(}\|p(a)-p(b)\|^{2}\big{)}_{(a,b)\in E}$. The graph $G$ is said to be generically rigid if $\Psi_{G}$ has finite fibers over general points of $\operatorname{Image}(\Psi_{G})$. Note that $\dim\operatorname{Image}(\Psi_{G})=2v-3$ if $G$ is generically rigid, and thus, if $|E|>2v-3$, there is some $E^{\prime}\subsetneq E$ such that $G^{\prime}=([v],E^{\prime})$ is generically rigid. Hence, a generically rigid graph $G$ is said to be minimally rigid if $\Psi_{G}$ is further dominant. By combining Theorem A with [LamanSphere], we obtain a new proof of Laman’s theorem [PollaczekGeiringer, Laman], stated below.
@@
A graph $G=([v],E)$ is minimally rigid if and only if it has $32v-$ 3 edges and every induced subgraph on $v^{\prime}>1$ vertices has at most $32v^{\prime}-$ 3 edges.
@@
Let $S=\{(x,y,z)\in\mathbb{R}^{3}:x^{2}+y^{2}+z^{2}=1\}$ be the 2-dimensional sphere, and consider the map $\widetilde{\Psi}_{G}\colon S^{[v]}/SO_{3}(\mathbb{R})\to\mathbb{R}^{E}$ given by $[p\colon[v]\to S]\mapsto\big{(}\frac{1-p(a)\bullet p(b)}{2}\big{)}_{(a,b)\in E}$, where $\bullet$ denotes the standard dot product, which is essentially the arc length between $p(a)$ and $p(b)$. A direct argument [Pogorelov] shows that a graph is minimally rigid (in the plane) if and only if it is minimally rigid in the sphere, i.e., the map $\widetilde{\Psi}_{G}$ is generically finite. By dimensional considerations, this happens only if $|E|=2v-3$, and in this case we need check whether $\widetilde{\Psi}_{G}$ is dominant, which can be done after base changing to $\mathbb{C}$. [LamanSphere] showed that the degree of the complexified map $\widetilde{\Psi}_{G,\mathbb{C}}$ is equal to a cross-ratio degree on $\overline{M}_{0,2v}$, as described in (iii). Translating the Cerberus condition on subsets of $[2v]$ of the form $\{a,b,a+v,b+v\}$ for $(a,b)\in E$ to a condition on induced subgraphs of $G$ and applying Theorem A yields the desired result. ∎
@@ 3. Recursion
We prepare with two lemmas about $X_{S,i}$ and boundary divisors on $\overline{M}_{0,n}$. For a partition $[n]=P\sqcup Q$ with $|P|\geq 2$ and $|Q|\geq 2$, let $D_{P|Q}$ be the corresponding boundary divisor, i.e.,
$D_{P|Q}=\text{the closure of }\left\{C\in\overline{M}_{0,n}:\begin{matrix}\text{$C$ has two $\mathbb{P}^{1}$-components, each with distinct marked points}\\ \text{labelled by $P$ or $Q$ respectively}\end{matrix}\right\}.$
@@
where the sum ranges over all bipartitions $[n]=P\sqcup Q$ such that $\{a,i\}\subseteq P$ and $S\setminus\{a,i\}\subseteq Q$.
@@
Consider the forgetful map $f_{S\setminus a}\colon\overline{M}_{0,S}\to\overline{M}_{0,S\setminus a}$. By [ArabelloCornalba, Lemma 3.1], we have that $\psi_{i}=X_{S\setminus a,i}+D_{\{a,i\}|(S\setminus\{a,i\})}$ on $\overline{M}_{0,S}$. The pullback of any $D_{P^{\prime}|Q^{\prime}}$ along $f_{S}\colon\overline{M}_{0,n}\to\overline{M}_{0,S}$ is the sum of the boundary divisors on $\overline{M}_{0,n}$ corresponding to the partitions obtained by adding elements of $[n]\setminus S$ to either $P^{\prime}$ or $Q^{\prime}$. The result then follows. ∎
@@
Let $(S,i)$ be a pair with $S\subseteq[n]$, $|S|\geq 4$, and $i\in S$. For a partition $[n]=P\sqcup Q$ with $|P|\geq 2$ and $|Q|\geq 2$, we have
@@
$P$$Q$$i$$P$$Q$$i$$P$$Q$$i$$j$
Figure 1. Illustrations of rational curves in the three cases of Lemma 3.2. The left rational component supports marked points indexed by $P$, the black round marked points are not indexed by $S$, and the blue hearts are marked points indexed by $S$.
@@
  • For $R\in\{P,Q\}$, the collection $\mathcal{S}_{R}:=\{\widehat{\operatorname{res}}_{R}(S_{j}):j\in\{2,\ldots,n-3\},|S_{j}\cap R|\geq 3\}$ has size $|R|-2$ and satisfies the Cerberus conditions on $R\cup\{*_{1}\}$, where we label * after $S_{1}$ to disambiguate.
@@
It thus suffices to exhibit a bijection between the set of good binary trees and the set of trivalent trees with the prescribed property. To go from a good binary tree to a trivalent tree, we first recursively construct the trivalent trees $\mathcal{T}_{P}$ and $\mathcal{T}_{Q}$ corresponding to $P\cup\{*\}$ and $Q\cup\{*\}$, respectively. We combine $\mathcal{T}_{P}$ and $\mathcal{T}_{Q}$ by identifying the edges connected to * in each tree such that the two *’s identify with distinct vertices. We label this edge by $*_{i}$, where $S_{i}$ is the set split at this node of the tree. At the leaves of the form $(D,\varnothing)$, where $|D|=3$, we connect the three vertices to a single internal vertex. Let $\mathcal{T}_{n}$ be the final trivalent tree.
@@
To see why $\mathcal{T}_{n}$ has the prescribed property, consider an arbitrary set $S_{j}$ with $j\in[n-3]$ and look at the internal node of the good binary tree in which $S_{i}$ is split into $\{a^{*}_{i},b^{*}_{i}\}$ and $\{c^{*}_{i},d^{*}_{i}\}$, where the *’s signify that the labels may have changed (repeatedly) during the recursion. The descents of $S_{i}$ only split $S_{j}$’s with $j>i$. Thus, when the edges corresponding to these larger $S_{j}$’s are contracted, we have that $a^{*}_{i}$ and $b^{*}_{i}$ are each adjacent to one side of $*_{i}$ while $c^{*}_{i}$ and $d^{*}_{i}$ are adjacent to the other side. As we traverse upward from the node splitting $S_{j}$ to the root of the tree, we will connect $a^{*}_{i}$ by a path (possibly of length 00) to $a_{i}$, and so forth, however none of these paths use $*_{i}$. Thus, the path from $a_{i}$ to $b_{i}$ and the path from $c_{i}$ to $d_{i}$ are connected by a single edge once $*_{i+1},\ldots,*_{n-3}$ are contracted, as desired.
@@
To invert the bijection, let $\mathcal{T}$ be a given trivalent tree. Label the internal edges of $\mathcal{T}$ as $*_{n_{3}},\ldots,*_{1}$ based on which edge is contracted when processing $S_{n-3},\ldots,S_{1}$, respectively. Replace $*_{1}$ with two disconnected vertices labeled *, each connected to one of the two internal vertices incident to $*_{1}$. Let $P$ and $Q$ be the subsets $[n]$ on each half of the split, with $a_{1},b_{1}\in P$; and let $\mathcal{T}_{P}$ and $\mathcal{T}_{Q}$ be the resulting trivalent trees. For each $j\in\{2,\ldots,n-3\}$, let $\mathcal{S}_{P}=\{\widehat{\operatorname{res}}_{P}(S_{j}):j\in\{2,\ldots,n-3\},|S_{j}\cap R|\geq 3\}$. Note that $\widehat{\operatorname{res}}_{P}(S_{j})\in\mathcal{S}_{P}$ if and only if $*_{j}$ is on the $P$-side of the tree split. From this, we can see that $\mathcal{S}_{P}$ satisfies the Cerberus condition: for any subset of set indices $J$, contract the internal edges of $\mathcal{T}_{P}$ not indexed by $J$ and delete any leaf whose label does not lie in $\bigcup_{j\in J}\widehat{\operatorname{res}}_{P}(S_{j})$. Each internal node has degree at least $3$ and there are $|J|$ internal edges, so there are at least $|J|+3$ leaves, as desired. Therefore, $P\sqcup Q$ is a good partition and we can repeate this procedure recursively. In the base case of trivalent trees with a single internal node, we map it to the label $(D,\varnothing)$, where $D$ is the set of labels of the three leaves. ∎
11112222333344445555666677771111222233334444555566667777

### v4

@@ 1. Introduction
For a finite set $S$ with cardinality $|S|\geq 3$, let $\overline{M}_{0,S}$ be the Deligne–Mumford–Knudsen moduli space of stable rational curves with marked points labeled by $S$. We write $\overline{M}_{0,n}$ for $\overline{M}_{0,[n]}$, where $n]\coloneqq\{1,\dots,n\}$ for an integer $n\geq 3$. For a pair $(S,i)$ consisting of a subset $S\subseteq[n]$ with $|S|\geq 3$ and an element $i\in S$, let
[…]
@@
For a collection $(S_{1},i_{1}),\dots,(S_{n-3},i_{n-3})$ of n3n-3 such pairs, we define its Kapranov degree as the intersection number
@@
[…]
@@
    When $|S_{j}|=4$ for all $j$, Kapranov degrees in this case were studied by Silversmith [Silversmith] under the name of cross-ratio degrees, because the divisor class $X_{S,i}$ when $|S|=4$ is the pullback of the hyperplane class under the map $\overline{M}_{0,n}\to\mathbb{P}^{1}$ given by taking the cross-ratio of the four points marked by $S$. That is, the Kapranov degree in this case can be interpreted as the number of ways to choose $n$ marked points on $\mathbb{P}^{1}$ with n3n-3 prescribed cross-ratios.
@@
    When further $n=2v$, and the size 4 subsets $S_{j}$ are obtained from a graph $G$ with vertex set $[v]$ and 2v32v-3 edges by assigning to each edge $(a,b)$ the subset $\{a,b,a+v,b+v\}$, the cross-ratio degree counts the number of (complex) realizations of $G$ on the $2$-dimensional sphere up to isometries for any fixed generic edge lengths of $G$ [LamanSphere].
@@
    When $i_{j}=n$ for all $j$, the Kapranov degrees were computed in [CTHypertrees, Theorem 3.2]. In this case, the Kapranov degree is either 0 or $1$.
@@
Our second result is a recursive formula for Kapranov degrees, described as follows. Let $\{\star\}$ be a singleton disjoint from $[n]$. Given a partition $n]=P\sqcup Q$ with $|P|\geq 2$ and $|Q|\geq 2$, and a pair $(S,i)$ with $S\subseteq[n]$, $|S|\geq 3$, and $i\in S$, we define an element $\operatorname{res}_{P}(S,i)$ in the cohomology of $\overline{M}_{0,P\cup\star}$ by
@@
An illustration of these cases in Figure 1 may be helpful for visualizing the above conditions. We similarly define $\operatorname{res}_{Q}(S,i)$ by replacing $P$ with $Q$ in the above. From the definition, it follows that $\operatorname{res}_{P}(S,i)=1$ if and only if $\operatorname{res}_{Q}(S,i)\neq 1$ and vice versa.
@@
When $|S_{j}|=4$ for all $j\in[n-3]$, i.e., the case of cross-ratio degrees, ˜B was proven in [LamanSphere]. It was also proven in [Goldner, Lemma 3.11] using tropical geometry (see [Silversmith]). We prove Theorem B in Section 3 by writing $X_{S,i}-X_{S\setminus a,i}$ as a sum of boundary divisors and using that the boundary divisors are products of smaller $\overline{M}_{0,n}$’s. Theorem B gives a useful algorithm for computing Kapranov degrees. [^1] An implementation in SageMath is available at https://github.com/MattLarson2399/KapranovDegrees. For the case of cross-ratio degrees, the recursion further admits a combinatorial interpretation in terms of trivalent trees (Corollary 3.3).
@@
Lastly, we give an explicit formula for the special case of Kapranov degrees where all the $S_{j}$ share a subset of size $3$, say $\{n-2,n-1,n\}$, and each $i_{j}$ lies in $\{n-2,n-1,n\}$. For a pair $(T,j)$ consisting of $T\subseteq[n-3]$ and $j\in\{n-2,n-1,n\}$, let $Y_{T,j}=X_{T\cup\{n-2,n-1,n\},j}$. For n3n-3 such pairs $(T_{1},j_{1}),\dotsc,(T_{n-3},j_{n-3})$, we define a $3$-transversal to be a map $t\colon[n-3]\to\{n-2,n-1,n\}$ such that there is a bijection $m\colon[n-3]\to[n-3]$ satisfying $m(i)\in T_{i}$ and $t(m(i))=j_{i}$ for all $i\in[n-3]$.
@@ 2.1. Reduction to the case of cross-ratio degrees
We will apply the Hall–Rado theorem to the matroid known as the third Dilworth truncation of the boolean matroid, denoted $D_{3}(U_{n,n})$. See, for example, [Whittle]. Its ground set is the collection $\binom{[n]}{4}$ of all subsets of $[n]$ of size $4$, and its set of bases consists of collections of n3n-3 subsets of size $4$ that satisfy the Cerberus condition. In [Athanasiadis, Theorem 3.2], the flats of $D_{3}(U_{n,n})$ are calculated to be as follows. They are in bijection with anti-chains $\{T_{1},\dotsc,T_{k}\}$ in the poset of subsets of $[n]$ that satisfy the conditions
@@
For the claim, we first note that the closure of $T_{S,i}$ in $D_{3}(U_{n,n})$ is the flat $F_{S}$ corresponding to $\{S\}$ because $F_{S}$ contains $T_{S,i}$ and they both have rank |S|3|S|-3. We now need show that the collection of flats $F_{S_{1}},\dotsc,F_{S_{n-3}}$ of $D_{3}(U_{n,n})$ satisfies (Hall–Rado). Suppose on the contrary that it fails (Hall–Rado), and take a minimal $J\subseteq[n-3]$ witnessing the failure. For each $j\in J$, because the flat $F_{S_{j}}$ is connected, it is contained in a connected component of the join $\bigvee_{j\in J}F_{S_{j}}$. Since ranks add over connected components, the minimality of $J$ implies that $\bigvee_{j\in J}F_{S_{j}}$ is a connected flat of rank less than $|J|$ in $D_{3}(U_{n,n})$. But the smallest connected flat containing $F_{S_{j}}$ for all $j\in J$ is the flat $F_{\bigcup_{j\in J}S_{j}}$, which has rank $|\bigcup_{j\in J}S_{j}|-3\geq|J|$ because of the Cerberus condition. ∎
@@ 2.2. Positivity for cross-ratio degrees
It suffices to show that the $(n-3)\times n$ matrix whose rows are given by the $d\operatorname{cr}_{S_{i}}$ has full rank; we may check this after scaling each row to clear denominators. By the computation above, if $S_{j}=\{j_{1},j_{2},j_{3},j_{4}\}$, then the $i$-th row has $(-1)^{\ell+1}\prod_{a<b\in S_{i},a,b\not=j_{\ell}}(x_{a}-x_{b})$ in the $j_{\ell}$-th entry and is zero otherwise.
Let $M$ be the $(n-3)\times n$ matrix whose $i$-th column is given by $(1,x_{i},\dotsc,x_{i}^{n-4})$. We apply Theorem 2.5 with m=n3m=n-3 and each $T_{i}=S_{i}$. Because $S_{1},\dotsc,S_{n-3}$ satisfies the Cerberus condition, there is an invertible $(n-3)\times(n-3)$ matrix $G$ such $(GM)_{ij}=0$ unless $j\in S_{i}$. By Proposition 2.6, we may pick $G$ such that $(GM)_{ij}=\prod_{k\in[n]\setminus S_{i}}(x_{j}-x_{k})$.
@@
Because the GM-MDS theorem holds over a field of any characteristic, Proposition 2.4 implies that, in any characteristic, the rational functions corresponding to cross-ratios form a separating transcendence basis when the Cerberus condition is satisfied. This implies that the map $\overline{M}_{0,n}\to(\mathbb{P}^{1})^{n-3}$ given by n3n-3 cross-ratios is generically étale when it is dominant.
@@
Let $I$ be the group of Euclidean isometries of the plane $\mathbb{R}^{2}$. For a graph $G$ with vertex set $[v]$ and edges $E$, consider the map $\Psi_{G}\colon(\mathbb{R}^{2})^{[v]}/I\to\mathbb{R}^{E}$ given by $p\colon[v]\to\mathbb{R}^{2}]\mapsto\big{(}\|p(a)-p(b)\|^{2}\big{)}_{(a,b)\in E}$. The graph $G$ is said to be generically rigid if $\Psi_{G}$ has finite fibers over general points of $\operatorname{Image}(\Psi_{G})$. Note that $\dim\operatorname{Image}(\Psi_{G})=2v-3$ if $G$ is generically rigid, and thus, if |E|>2v3|E|>2v-3, there is some $E^{\prime}\subsetneq E$ such that $G^{\prime}=([v],E^{\prime})$ is generically rigid. Hence, a generically rigid graph $G$ is said to be minimally rigid if $\Psi_{G}$ is further dominant. By combining Theorem A with [LamanSphere], we obtain a new proof of Laman’s theorem [PollaczekGeiringer, Laman], stated below.
@@
A graph $G=([v],E)$ is minimally rigid if and only if it has 2v32v-3 edges and every induced subgraph on $v^{\prime}>1$ vertices has at most $2v^{\prime}-3$ edges.
@@
Let $S=\{(x,y,z)\in\mathbb{R}^{3}:x^{2}+y^{2}+z^{2}=1\}$ be the 2-dimensional sphere, and consider the map $\widetilde{\Psi}_{G}\colon S^{[v]}/SO_{3}(\mathbb{R})\to\mathbb{R}^{E}$ given by $p\colon[v]\to S]\mapsto\big{(}\frac{1-p(a)\bullet p(b)}{2}\big{)}_{(a,b)\in E}$, where $\bullet$ denotes the standard dot product, which is essentially the arc length between $p(a)$ and $p(b)$. A direct argument [Pogorelov] shows that a graph is minimally rigid (in the plane) if and only if it is minimally rigid in the sphere, i.e., the map $\widetilde{\Psi}_{G}$ is generically finite. By dimensional considerations, this happens only if |E|=2v3|E|=2v-3, and in this case we need check whether $\widetilde{\Psi}_{G}$ is dominant, which can be done after base changing to $\mathbb{C}$. [LamanSphere] showed that the degree of the complexified map $\widetilde{\Psi}_{G,\mathbb{C}}$ is equal to a cross-ratio degree on $\overline{M}_{0,2v}$, as described in (iii). Translating the Cerberus condition on subsets of $[2v]$ of the form $\{a,b,a+v,b+v\}$ for $(a,b)\in E$ to a condition on induced subgraphs of $G$ and applying Theorem A yields the desired result. ∎
@@ 3. Recursion
We prepare with two lemmas about $X_{S,i}$ and boundary divisors on $\overline{M}_{0,n}$. For a partition $n]=P\sqcup Q$ with $|P|\geq 2$ and $|Q|\geq 2$, let $D_{P|Q}$ be the corresponding boundary divisor, i.e.,
@@
With $\{\star\}$ a singleton disjoint from $[n]$, the gluing map $\operatorname{gl}_{P|Q}\colon\overline{M}_{0,P\cup\star}\times\overline{M}_{0,Q\cup\star}\to\overline{M}_{0,n}$, which glues two marked curves $C_{1}\in\overline{M}_{0,P\cup\star}$ and $C_{2}\in\overline{M}_{0,Q\cup\star}$ by identifying the points marked by $\star$, is an isomorphism onto the boundary divisor $D_{P|Q}$. Let $\pi_{P}\colon\overline{M}_{0,P\cup\star}\times\overline{M}_{0,Q\cup\star}\to\overline{M}_{0,P\cup\star}$ be the projection onto the first factor, and similarly for $\pi_{Q}$.
@@
where the sum ranges over all bipartitions $n]=P\sqcup Q$ such that $\{a,i\}\subseteq P$ and $S\setminus\{a,i\}\subseteq Q$.
@@
Consider the forgetful map $f_{S\setminus a}\colon\overline{M}_{0,S}\to\overline{M}_{0,S\setminus a}$. By [ArabelloCornalba, Lemma 3.1], we have that $\psi_{i}=X_{S\setminus a,i}+D_{\{a,i\}|(S\setminus\{a,i\})}$ on $\overline{M}_{0,S}$. The pullback of any $D_{P^{\prime}|Q^{\prime}}$ along $f_{S}\colon\overline{M}_{0,n}\to\overline{M}_{0,S}$ is the sum of the boundary divisors on $\overline{M}_{0,n}$ corresponding to the partitions obtained by adding elements of $n]\setminus S$ to either $P^{\prime}$ or $Q^{\prime}$. The result then follows. ∎
@@
Let $(S,i)$ be a pair with $S\subseteq[n]$, $|S|\geq 4$, and $i\in S$. For a partition $n]=P\sqcup Q$ with $|P|\geq 2$ and $|Q|\geq 2$, we have
@@
PPQQiiPPQQiiPPQQiijj
Figure 1. Illustrations of rational curves in the three cases of Lemma˜3.2. The left rational component supports marked points indexed by $P$, the black round marked points are not indexed by $S$, and the blue hearts are marked points indexed by $S$.
@@
  • For $R\in\{P,Q\}$, the collection $\mathcal{S}_{R}:=\{\widehat{\operatorname{res}}_{R}(S_{j}):j\in\{2,\ldots,n-3\},|S_{j}\cap R|\geq 3\}$ has size |R|2|R|-2 and satisfies the Cerberus conditions on $R\cup\{*_{1}\}$, where we label * after $S_{1}$ to disambiguate.
@@
It thus suffices to exhibit a bijection between the set of good binary trees and the set of trivalent trees with the prescribed property. To go from a good binary tree to a trivalent tree, we first recursively construct the trivalent trees $\mathcal{T}_{P}$ and $\mathcal{T}_{Q}$ corresponding to $P\cup\{*\}$ and $Q\cup\{*\}$, respectively. We combine $\mathcal{T}_{P}$ and $\mathcal{T}_{Q}$ by identifying the edges connected to * in each tree such that the two *’s identify with distinct vertices. We label this edge by $_{i}$, where $S_{i}$ is the set split at this node of the tree. At the leaves of the form $(D,\varnothing)$, where $|D|=3$, we connect the three vertices to a single internal vertex. Let $\mathcal{T}_{n}$ be the final trivalent tree.
@@
To see why $\mathcal{T}_{n}$ has the prescribed property, consider an arbitrary set $S_{j}$ with $j\in[n-3]$ and look at the internal node of the good binary tree in which $S_{i}$ is split into $\{a^{*}_{i},b^{*}_{i}\}$ and $\{c^{*}_{i},d^{*}_{i}\}$, where the *’s signify that the labels may have changed (repeatedly) during the recursion. The descents of $S_{i}$ only split $S_{j}$’s with $j>i$. Thus, when the edges corresponding to these larger $S_{j}$’s are contracted, we have that $a^{*}_{i}$ and $b^{*}_{i}$ are each adjacent to one side of $_{i}$ while $c^{*}_{i}$ and $d^{*}_{i}$ are adjacent to the other side. As we traverse upward from the node splitting $S_{j}$ to the root of the tree, we will connect $a^{*}_{i}$ by a path (possibly of length 0) to $a_{i}$, and so forth, however none of these paths use $_{i}$. Thus, the path from $a_{i}$ to $b_{i}$ and the path from $c_{i}$ to $d_{i}$ are connected by a single edge once $_{i+1},\ldots,*_{n-3}$ are contracted, as desired.
@@
To invert the bijection, let $\mathcal{T}$ be a given trivalent tree. Label the internal edges of $\mathcal{T}$ as $_{n_{3}},\ldots,*_{1}$ based on which edge is contracted when processing $S_{n-3},\ldots,S_{1}$, respectively. Replace $_{1}$ with two disconnected vertices labeled *, each connected to one of the two internal vertices incident to $_{1}$. Let $P$ and $Q$ be the subsets $[n]$ on each half of the split, with $a_{1},b_{1}\in P$; and let $\mathcal{T}_{P}$ and $\mathcal{T}_{Q}$ be the resulting trivalent trees. For each $j\in\{2,\ldots,n-3\}$, let $\mathcal{S}_{P}=\{\widehat{\operatorname{res}}_{P}(S_{j}):j\in\{2,\ldots,n-3\},|S_{j}\cap R|\geq 3\}$. Note that $\widehat{\operatorname{res}}_{P}(S_{j})\in\mathcal{S}_{P}$ if and only if $_{j}$ is on the $P$-side of the tree split. From this, we can see that $\mathcal{S}_{P}$ satisfies the Cerberus condition: for any subset of set indices $J$, contract the internal edges of $\mathcal{T}_{P}$ not indexed by $J$ and delete any leaf whose label does not lie in $\bigcup_{j\in J}\widehat{\operatorname{res}}_{P}(S_{j})$. Each internal node has degree at least $3$ and there are $|J|$ internal edges, so there are at least $|J|+3$ leaves, as desired. Therefore, $P\sqcup Q$ is a good partition and we can repeate this procedure recursively. In the base case of trivalent trees with a single internal node, we map it to the label $(D,\varnothing)$, where $D$ is the set of labels of the three leaves. ∎
$11223344556677$